OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10

OpenStudy (math&ing001):

We isolate the x in the second equation by adding 2y on both sides then multiplying by (-1).

OpenStudy (math&ing001):

Can you do that @kaylamarie081 ? I'll walk you through if you can't.

OpenStudy (anonymous):

I added 2y to both sides on the second equation and I got -x= 30. It doesn't look right though. Help me please

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

I don't know where you got that 30 but this is the step by step : -x - 2y = 10 + 2y +2y -x = 10 +2y -x*(-1) = (10+2y)*(-1) x= -10 - 2y So the expression you'll substitute into the first equation is (-10-2y).

OpenStudy (anonymous):

Can you help me with another one please

OpenStudy (anonymous):

Solve the following system of equations. x - y = 16 2x + y = 2

OpenStudy (math&ing001):

Ok but I will not solve it for you. You'll have to do the work.

OpenStudy (math&ing001):

Isolate x in the first equation. It's just like you're solving for x.

OpenStudy (math&ing001):

@kaylamarie081

OpenStudy (anonymous):

OK, just a minute. I'm trying to solve it on paper.

OpenStudy (math&ing001):

Alright, take your time ^^

OpenStudy (anonymous):

Am I doing it right so far

OpenStudy (anonymous):

I get confused on these kinda equations.

OpenStudy (math&ing001):

Isolating x should give you x=16-y

OpenStudy (anonymous):

Okay so I messed up at 2(16)+y=2?

OpenStudy (math&ing001):

Sorry I took so long, I was away :)

OpenStudy (anonymous):

That's fine. :)

OpenStudy (math&ing001):

Actually you messed up on the first equation, and that's what got the second one wrong. If you replace the x with (16-y) on the second you should be fine.

OpenStudy (anonymous):

:(

OpenStudy (math&ing001):

Where are you stuck ?

OpenStudy (anonymous):

Confused about everything.. K so I have to solve that?^^^

OpenStudy (math&ing001):

Yep !

OpenStudy (math&ing001):

It IS right so far, go on =3

OpenStudy (anonymous):

It equals 30?

OpenStudy (math&ing001):

Yeah !

OpenStudy (anonymous):

You're awesome. Thank you so much. Defintley couldn't of done it without you!(:

OpenStudy (math&ing001):

Now we replace it on the first equation to find x. x=16-y

OpenStudy (anonymous):

K

OpenStudy (anonymous):

Wait nvm. I need to find x not y

OpenStudy (math&ing001):

Exactly ! Because you already found y.

OpenStudy (math&ing001):

You went too far, all you have to do is replace y by its value in here : x = 16-y = 16-30 = -14

OpenStudy (math&ing001):

In case you're still confused, here's what we did so far :

OpenStudy (anonymous):

Thanks.

OpenStudy (math&ing001):

Welcome !

OpenStudy (anonymous):

Could you help me on this word problem?

OpenStudy (math&ing001):

Well the distance doesn't change, so that's how you'll find the speed on still water. Distance = 45 * 3 = speedonstillwater * 2 So speedonstillwater = 45*3/2

OpenStudy (anonymous):